OpenStudy (anonymous):

Find all real solutions of the equation. (Enter your answers as a comma-separated list. If there is no real solution, enter NO REAL SOLUTION.) x − (_/8 − 5x) = 0

OpenStudy (anonymous):

x - sqrt8 - 5x = 0 -4x - sqrt8 = 0 -4x = sqrt8 -x = sqrt8/4 gives the answer x = - 0.7071 or + 0.7071